### v1

# Radio-loudness statistics of quasars from Quaia–VLASS

Abstract
Context. Quasars are objects of high interest for applications in extragalactic astrophysics, cosmology, and astrometry. One of their useful qualities is owed to their relativistic jets: they can be radio-loud. However, the fraction of radio-loud vs. radio-quiet quasars is subject to ongoing investigations, where the statistical power is limited by the low number of known quasars with radio counterparts.
Aims. In this analysis, we revisited the radio-loudness statistics of quasars by significantly expanding the pool of known sources. Our main goal was to create a new, value-added catalogue of quasars with information about their extinction-corrected magnitudes, radio flux density, possible contamination levels, and other data flags, besides their sky coordinates and photometric redshifts.
Methods. We cross-matched the optical Quaia catalogue of about $1.3$ million quasars (selected from the Gaia data set) with $1.9$ million sources from the Very Large Array Sky Survey (VLASS) radio catalogue. We explored different thresholds for the matching radius, balancing the completeness and purity of the resulting Quaia–VLASS catalogue, and found that $1.5\arcsec$ is a sufficient choice.
Results. Our main finding is that the radio-loud fraction of quasars is in good agreement with previous works ($<10\%$), and there is no significant large-scale pattern in radio-loudness across the sky. The exact estimate depends on the $G$-band magnitude limit, and we observed some weak trends with redshift and absolute optical magnitude, possibly indicating remnant systematic effects in our data sets. The cross-matched Quaia–VLASS catalogue with 43,650 sources is available to the public for future analyses.
Conclusions. This latest census of QSOs with radio counterparts will facilitate further investigations of the dichotomy of radio-loud and radio-quiet quasars, and it may also support other lines of statistical investigations using quasars in cosmology and astrophysics.
Key Words.:
catalogs – surveys – galaxies: active – quasars: general – radio continuum: galaxies – large-scale structure of Universe
1 Introduction
Active galactic nuclei (AGN) are among the most powerful and dynamic objects in the Universe (e.g. Padovani et al., 2017), usually outshining their host galaxies. They can be found over a wide range of distances from the local Universe up to extremely high redshifts ($z\gtrsim 10$, Goulding et al., 2023). AGN play an important role in astrophysics, astrometry, and cosmology. They are located at the centres of galaxies and powered by accretion onto supermassive black holes (SMBHs) which range from millions to billions of solar masses. As material spirals into the SMBH, it heats up and emits huge amounts of energy across the entire electromagnetic spectrum, from radio to $\gamma$-rays.
Members of a small but spectacular subset of AGN exhibit intense radio emission originating from synchrotron radiation of relativistic charged particles moving in magnetized plasma jets launched from the close vicinity of SMBHs (e.g. Urry & Padovani, 1995; Blandford et al., 2019). Studying these so-called radio-loud, or rather jetted (Padovani, 2017) AGN is crucial for several reasons. Their jets are among the most extreme particle accelerators in the Universe, giving unique insight into relativistic effects, strong magnetic fields, and high-energy particle interactions. The radiative and mechanical feedback of AGN jets can influence the evolution of their host galaxies by regulating star formation (e.g. Fabian, 2012; Hardcastle & Croston, 2020).
Figure 1: The distribution of the $1.1$ million Quaia sources that fall within the joint coverage with the VLASS survey ($\delta>-40\degr$). The dust obscuration in the Milky Way disc results in the detection of fewer QSOs in the optical.
Owing to their presence in the widest range of cosmological distances, AGN are excellent probes of the early Universe and offer the possibility of contributing to the studies of cosmological models (e.g. Vermeulen & Cohen, 1994; Gurvits et al., 1999). AGN also trace cosmic structures and thus are important for understanding the large-scale distribution of matter (see e.g. Kovács et al., 2022; Piccirilli et al., 2024). In astrometry, bright radio-emitting AGN define the most precise reference frame through high-resolution very long baseline interferometry (VLBI) observations (Charlot et al., 2020).
A subclass of AGN, quasars (originally meaning quasi-stellar radio sources) were discovered as distant extragalactic objects in the early 1960’s (Schmidt, 1963). Not much later, it turned out that the majority of the so-called quasi-stellar objects (QSOs, also called quasars for simplicity) are not strong radio emitters (Sandage, 1965). Nowadays, the number of known quasars with measured spectroscopic redshifts exceeds 1 million (Flesch, 2023). The fraction of radio-loud quasars among the whole population is known to be $\lesssim 10\%$ (e.g. Ivezić et al., 2002; Kellermann et al., 2016).
There are various working definitions of radio loudness in the literature. One can use the radio-to-optical flux density ratios as originally proposed by Schmidt (1970). Perhaps the most widely used definition is based on the ratio of the rest-frame $5$-GHz radio and 4400-Å optical flux densities (Kellermann et al., 1989). If this ratio $R$ exceeds 10 then the object is considered radio-loud. Sramek & Weedman (1980) used 2500-Å optical flux densities. Given the well-known flux density variability of quasars and the fact that optical and radio measurements are rarely contemporaneous, $R$ values themselves can be uncertain.
The above two definitions are broadly consistent with each other for the quasar class as a whole (see a recent example in Krezinger et al., 2024). Alternatively, dictated by practical aspects such as the availability of measurements, it is customary to use other radio and optical bands, e.g. $1.4$ GHz in the radio and the $i$-band in the optical (like in Ivezić et al., 2002; Baloković et al., 2012). Note that Terashima & Wilson (2003) defined radio loudness as the ratio of radio ($5$ GHz) and X-ray ($2-10$ keV band) luminosities that can be applied for heavily obscured AGN. A fundamentally different approach is defining radio loudness by setting a threshold based on the (monochromatic) radio luminosity (e.g. Peacock et al., 1986; Miller et al., 1990). Interestingly, the two types of definitions provide generally similar classifications, as discussed in Appendix C of Ivezić et al. (2002).
Figure 2: The distribution of the $1.9$ million VLASS2 QL sources after applying quality flags (see Table 1). Unlike for optical QSOs, the radio coverage is virtually unaffected by the Galactic disc.
The ‘classical’ definition of radio loudness based on radio and optical flux density ratios is the best applicable for quasars where the optical emission is dominantly related to the accretion disk while the radio emission comes mainly from the jets (Padovani, 2017). Most early publications found a bimodal distribution of radio-loud and radio-quiet (in fact, not totally free from radio emission, rather radio-weak) quasars (e.g. Strittmatter et al., 1980; Kellermann et al., 1989; Miller et al., 1990; Ivezić et al., 2002). While the debate about this dichotomy is not settled yet, a consensus seems to emerge that radio loudness has a continuous distribution with no sharp gap between radio-loud and radio-quiet sources (see a review in Kellermann et al., 2016).
For a long time, studies of radio loudness were naturally limited to small samples of relatively bright AGN. The emergence of sensitive large radio and optical sky surveys (e.g. Ivezić et al., 2002; Gürkan et al., 2019; Arnaudova et al., 2024) made substantially larger samples available for investigations. Here we revisit the problem with modern optical and radio surveys, and perform various statistical tests of radio-loudness of quasars.
The paper is organized as follows. In Sect. 2, we introduce our observational data sets and the analysis methodology. Then, Sect. 3 contains our observational results, followed by a summary of our main conclusions in Sect. 4.
2 Data sets and methodology
The main goal of this work is to cross-match a large sample of optically detected quasars with a catalogue of radio sources, mainly following the footsteps of Ivezić et al. (2002), concentrating on the radio-loudness statistics. In this section, we present the input data sets and the technical details of the cross-matching process.
For the optical quasar catalogue, we chose the recently compiled Quaia (Storey-Fisher et al., 2024) because of its large sky coverage, high completeness, and high astrometric accuracy. For the radio data, we relied on the Very Large Array Sky Survey (VLASS, Lacy et al., 2020) which is superior to all previous large sky surveys in terms of positional accuracy, angular resolution, and sensitivity.
2.1 Quasar catalogue
For our cross-matched sample, we used the Quaia quasar catalogue (Storey-Fisher et al., 2024), culled from the Gaia space telescope’s data set (Gaia Collaboration et al., 2016). The QSOs of the Quaia catalogue are distinguished from the other types of sources by cross-matching the Gaia Data Release 3 (DR3) QSO candidates (Gaia Collaboration et al., 2023b, a; Delchambre et al., 2023) with infrared sources from the Wide-field Infrared Survey Explorer (WISE, Lang, 2014), applying colour cuts within the GaiaWISE photometry parameter space and cutting out sources with high proper motions.
More precisely, we used the Quaia catalogue restricted in magnitude to $G<20.5$, containing about $1.3$ million QSOs, which yields high completeness of quasars within the Gaia sky coverage. There is also a conservative Quaia catalogue, a subsample of the larger one, with a constraint of $G<20.0$, with the benefit of the latter constraint yielding a purer QSO sample with a higher fidelity of the sources being QSOs. We decided to use the more complete catalogue with $G<20.5$, because its contaminants are expected to be mostly stars and should have negligible emission in the radio.
Another Quaia data product that we used is the so-called selection function, which indicates the QSO completeness due to various systematic effects (see Storey-Fisher et al., 2024, for details). The sky coverage of the Quaia catalogue that we consider in this analysis is displayed in Fig. 1.
Figure 3: Left: the number of cross-matched sources as a function of a search radius parameter in arcseconds. We performed a robustness test using randomly perturbed sky coordinates, and found that a search radius of $1.5\arcsec$ results in a low level of contamination ($<0.1\%$) from false matches (right). We thus choose this threshold for our cross-matched Quaia–VLASS sample.
2.2 Radio sources
The other data set that we used is based on the ongoing VLASS (Lacy et al., 2020) at 3 GHz (S-band), conducted with the Very Large Array in New Mexico, in particular the VLASS Epoch 2 Quick Look (VLASS2 QL) catalogue. Unlike Quaia, the VLASS2 QL catalogue does not cover the full sky, and its 3.0 million radio sources are distributed at declinations $\delta>-40\degr$.
Further, VLASS is conducted in a single band, making it more difficult to deduct the class of the detected objects than for optical surveys with multiple bands, e.g. by segmentation in colour–colour space. Hence, cross-matching VLASS with another large-scale catalogue like Quaia is a straightforward choice for sampling radio-loud extragalactic objects. The sky area density of the VLASS2 QL catalogue, with applied quality flags, is displayed in Fig. 2.
Figure 4: The distribution of angular separation values when cross-matching the Quaia and VLASS samples, up to the $1.5\arcsec$ search radius threshold that we applied. The inset shows the closest matches below $\Delta\theta=0.2\arcsec$.
For our purposes, we applied the VLASS data flags explicitly recommended by the user manual (Gordon et al., 2023) to restrict ourselves to the most reliable data. We used the VLASS catalogue flag setting Duplicate_flag¡2 to get rid of any duplicated radio sources, originating from overlaps of the sky patches observed in the survey.
[…]
Table 1: Population statistics for the different data sets used.
Source catalogues $N_{\mathrm{sources}}$ Radio-loud
Quaia, full ($G<20.5$) 1,295,502 -
Quaia, $\delta>-40^{\circ}$ cut 1,084,545 -
VLASS2 QL 2,995,025 -
VLASS2 QL, flags applied 1,873,524 -
Quaia–VLASS (all) 43,650 $3.83\%\pm 0.02\%$
$18.514.0<G<$ 18.5 cut 9,395 $8.04\%\pm 0.09\%$
$19.518.5\leq G<$ 19.5 cut 15,763 $4.71\%\pm 0.04\%$
$20.519.5\leq G<$ 20.5 cut 18,492 $2.80\%\pm 0.02\%$
 [^1] Top: the number of sources contained in the data sets that we used. Bottom: The radio-loud fraction of sources in the full Quaia–VLASS cross-match and in subsets due to different cuts in $G$ optical magnitude (errors are based on Poisson statistics).
2.3 Cross-matching
In order to create a cross-matched sample, we determined the common, physically identical sources within Quaia and VLASS. We first identified for each Quaia source its closest counterpart in VLASS and then calculated the corresponding angular separation between their sky coordinates ($\Delta\theta$). This method has been previously used, among others, by Ivezić et al. (2002); Orosz & Frey (2013).
Figure 5: The extent of our Quaia–VLASS sample along the $G$ (optical) and $t$ (radio) bands. Radio-loud and radio-quiet sources are separated by the black line. We defined three bins in the $G$ band, which we later analyse separately in the rest of the paper when looking for brightness-dependent trends in radio loudness.
Next, we determined an optimal angular separation threshold $\Delta\theta_{\mathrm{t}}$, for which one can safely assume that sources with $\Delta\theta<\Delta\theta_{\mathrm{t}}$ are physically related in Quaia and VLASS, while sources with $\Delta\theta\geq\Delta\theta_{\mathrm{t}}$ are plausibly just chance alignments of non-identical sources. An optimal choice of $\Delta\theta_{\mathrm{t}}$ should be large enough to allow for a large number of proper physical cross-matches to be identified as such (high completeness of the sample). Simultaneously, $\Delta\theta_{\mathrm{t}}$ should be reasonably small to prevent contamination of the cross-matched population with chance-aligned sources (high-purity sample).
In Fig. 3, we present our findings in the context of the fraction of possibly false matches. In the left panel, $N_{\mathrm{cross}}$ is the cumulative number of cross-matches between Quaia and VLASS as a function of the coordinate separation up to the corresponding $\Delta\theta$ value. We also display $N_{\rm false}$, calculated with a method used previously by Orosz & Frey (2013). $N_{\rm false}$ measures a cross-matched fraction between VLASS and Quaia*, where Quaia* is a surrogate of the Quaia QSO catalogue with all sources rotated randomly by the same angle in right ascension. This ensures that $N_{\rm false}$ from cross-matching Quaia* and VLASS only contains unphysical chance alignments. Moreover, $N_{\rm false}$ should approximate the chance alignments between Quaia and VLASS very well, because Quaia* has the same source density as Quaia. The curve of $N_{\mathrm{corrected}}=N_{\mathrm{cross}}-N_{\rm false}$ demonstrates a saturation in the number of physical cross-matches at $\Delta\theta\approx 1.5\arcsec$ (displayed as a grey band).
Using the above estimate of the contamination, in Fig. 3 (right panel) we show the fraction of chance alignments as a function of $\Delta\theta$, which can be expressed as
$f(\theta)=\frac{N_{\rm false}(\theta)}{N_{\rm corrected}(\theta)}=\frac{N_{\rm false}(\theta)}{N_{\rm cross}(\theta)-N_{\rm false}(\theta)}.$ (1)
At $\Delta\theta\approx 1.5\arcsec$, we estimated $N_{\rm false}/N_{\rm corrected}\approx 0.1\%$ contamination, i.e. excellent purity. We thus decided to apply this $\Delta\theta_{\mathrm{t}}<1.5\arcsec$ threshold for our cross-matching procedure. We note that Ivezić et al. (2002) also used the same threshold for their analysis, which makes comparisons more straightforward.
Figure 4 shows a histogram of the Quaia–VLASS angular separations up to $\Delta\theta=1.5\arcsec$, indicating that most cross-matches are identified with $\Delta\theta<0.5\arcsec$ separation, with a peak at about $\Delta\theta\approx 0.07\arcsec$ (see inset of Fig. 4). We thus conclude that the construction of our cross-matched sample guarantees high completeness.
2.4 Reddening correction
As provided by Storey-Fisher et al. (2024), the $G$-band magnitudes of the Quaia sources are not corrected for Galactic extinction. We thus corrected the $G$-band data following the formula
$G=G^{\prime}-r\cdot\langle R_{\zeta}\rangle,$ (2)
where $G^{\prime}$ is the uncorrected magnitude, $G$ is extinction-corrected magnitude, and $r$ is the reddening in the given region of the sky as provided by Schlegel et al. (1998). Further, $\langle R_{\zeta}\rangle\approx 2.74$ is the extinction coefficient for the Gaia $G$-band as provided by Casagrande & VandenBerg (2018). Except stated otherwise, all subsequent references in this paper to the $G$-band magnitude refer to the corrected value from Eq. 2.
3 Results
Using the cross-matched Quaia–VLASS sources, we estimated the radio loudness of our sample in various subsets and using different cuts. We used the definition by Ivezić et al. (2002) for the radio magnitude $t$ as
$t=-2.5\log\left(\frac{F_{\mathrm{tot}}}{3631\mathrm{Jy}}\right),$ (3)
where $F_{\mathrm{tot}}$ is the total flux density of the source (in the $S$-band). In Fig. 5, all the cross-matched Quaia–VLASS sources are displayed in a $G-t$ diagram, including a binning along the $G$ axis for future tests. Again following Ivezić et al. (2002), we used the $G-t$ magnitude difference as a measure for the radio loudness of the sources. We defined the following two cases:
$\begin{array}[]{ll}G-t>2.5:&\text{{\bf radio-loud sources}}\\ G-t\leq 2.5:&\text{radio-quiet sources}\end{array}$ (4)
Table 2: Differences between our work and Ivezić et al. (2002).
$\hline\cr\hline\cr Present work&\text{Previous work}\\ \hline\cr\emph{Quaia}--VLASS&\text{SDSS--FIRST}\\ $G$-band + $S$-band&\text{$i$-band + $L$-band}\\ 43,650 cross-matches&\text{$3066$ cross-matches}\\ \hline\cr\end{array}$
This definition was derived from the equivalent calculation in Ivezić et al. (2002), $R_{m}=0.4(m-t)$, where $m$ is a magnitude in the optical, $t$ is the radio magnitude at $1.4$ GHz (in that example, $R_{m}>1$ corresponds to radio-loud and $R_{m}\leq 1$ to radio-quiet objects).
Figure 6: The radio-loud fraction $N_{\rm loud}/N_{\rm total}$ of QSOs in different bins as a function of $G$ (black curve) and the number $N_{\rm total}$ of Quaia quasars in these bins (blue curve). The colour-coding along $G$ is equivalent to the colour-coding in Fig. 5. The black curve shows that even though it is expected to underestimate the radio-loud fraction at $G>18.5$, such a trend of a sudden drop in the curve is not observed.
3.1 A high-purity catalogue of radio-loud quasars
In Fig. 5, a solid line depicts the border between radio-loud and radio-quiet sources, as expressed in Eq. 4. Sources are coloured according to bins in $G$, which correspond to optically bright, moderate, and faint bins (the number of sources within different $G$-bins is given in Table 1).
The choice of $G$-band range values for these bins is motivated by a desire to keep these splits simple, yet offering meaningful comparisons to Ivezić et al. (2002). We note that while we kept $t$ and radio loudness consistent for the sake of comparison, there are some inevitable technical differences as displayed in Table 2. The main addition is a significantly larger sky coverage, a larger cross-matched sample, and thus a reduction of possible biases in the analysis.
While the actual definitions of radio loudness may differ (as well as their range of validity), we note that of the 43,650 cross-matched Quaia–VLASS quasars shown in Fig. 5 the vast majority is clearly visible in the radio (41,496 of them are radio-loud, given our definition). Therefore, an important outcome of our work is a high-purity catalogue of quasars with radio counterparts, which may be used in various other projects in the future.
Considering the completeness of our radio-loud catalogue (besides its purity), one should also consider the sensitivity limits of the input optical and radio data sets, that may bias the radio-loudness statistics for faint sources. We provide various tests of these systematic effects in the following sections, applying different binning schemes and cuts to the Quaia–VLASS sample to estimate possible variations in the radio-loud fraction.
3.2 Radio-loud fraction vs. $G$ magnitude
When estimating the radio-loud fraction ($N_{\rm loud}/N_{\rm total}$) of Quaia–VLASS sources, it is reasonable to restrict the sample to a range where the sensitivity of the instruments is reliable, i.e. where the measured $N_{\rm loud}$ and $N_{\rm total}$ source counts are complete within a chosen magnitude limit.
Ivezić et al. (2002) applied such a restriction in the form of a cut in the optical $i$-band, providing a fiducial value for the fraction of radio-loud QSOs with $i<18.5$. This particular cut follows from their radio-loudness formula ($G-t>2.5$) and the sensitivity limit they work with in the radio ($t<16.0$, that is slightly more conservative than the nominal $t<16.5$ where they still find radio sources).
Figure 7: The $G-t$ magnitude difference, i.e. our proxy for radio loudness, is averaged in various bins. Top: we bin our catalogue on the Quaia selection function (completeness) values, and also illustrate the role of the Galactic extinction correction. Bottom: we bin our sample on redshift, and also on $G$ magnitudes (see colours), and identify weak trends at the lowest redshifts $(z<0.5)$.
In our work, we used the same Eq. 4 for defining radio loudness, and the VLASS has a similar sensitivity limit in $t$ (see Fig. 5). While this condition motivates a cut in the Quaia data at $G<18.5$, we tested the whole $G<20.5$ magnitude range to measure $N_{\rm loud}/N_{\rm total}$ for a comprehensive analysis.
Figure 5 depicts the origin of this unreliability. For fainter Quaia magnitudes, the radio magnitudes may fall outside of the sensitivity limit $t\approx 16.0$, which suggests lower and lower completeness when one intends to estimate the radio-loud fraction on the faint end (while the purity of the sample is expected to remain high, as we argued above).
In Fig. 6, we show how the radio-loud fraction of QSOs drops for fainter optical sources (which are naturally more abundant in the sample). We made the following observations:
  • The black line represents the radio-loud fraction of sources as a function of the $G$ magnitude, showing a linear drop in the radio-loudness fraction.
  • Also, we observed that the radio-loud fraction is not constant even in the range of the brightest sources (e.g. $G<17$), where high completeness is expected in both $N_{\rm loud}$ and $N_{\rm total}$.
  • Further, constraining the measurement of the radio-loud fraction to $G<18.5$ would ignore the bulk of the Quaia QSOs (see the blue curve in 6) and the majority of the cross-matched sources (see Table 1).
For the above three reasons, we decided to express the radio-loud fraction of the Quaia–VLASS sources as a function of $G$ up to $G=20.5$, and let future users of our catalogue decide how they wish to balance purity and completeness.
Figure 8: Radio-loud QSO fraction values in different parts of the sky are shown. We find significant fluctuations near the Milky Way, i.e. where the Quaia catalogue is knowingly less complete (although on average there are no clear biases near the Galactic plane either, see top panel of Fig. 7). In the cleanest areas of the sky, we do not identify any relevant anisotropy in the radio-loud fraction.
Nevertheless, if we apply the measure for the radio-loud fraction by Ivezić et al. (2002), then we arrive at a value of $N_{\rm loud}/N_{\rm total}\approx 8.04\%\pm 0.09\%$ radio-loud fraction for $G<18.5$, while they estimate $N_{\rm loud}/N_{\rm total}\approx 8\%\pm 1\%$ for a similarly bright, conservatively selected subset of their data ($i<18.5$). Again, we note the differences between the actual bands used by Ivezić et al. (2002) and our analysis, but we conclude that consistent radio-loud fraction values are observed for the brightest QSOs.
3.3 Radio loudness vs. QSO completeness
In Fig. 7, we show multiple tests that we performed by binning and modifying our data. In the context of the Quaia selection function (a measure for the completeness of QSO detections per pixel), we show on the top panel the role of the $G$-band extinction correction to $G-t$ (our proxy for the radio loudness), and the overall spread of the Quaia–VLASS catalogue as a function of this systematic parameter of the input data.
The colour of the data points corresponds to their $G$-band magnitudes (as shown in Fig. 5). Empty circles mark the average value of $G^{\prime}-t$ within a given bin of the selection function values, considering the raw $G^{\prime}$ magnitudes (not corrected for extinction). In contrast, full black circles show the binned average $G-t$ values after extinction correction.
From the comparison to the overall mean $G-t$ value of the whole Quaia–VLASS sample (horizontal line), we found that the extinction correction is effective in reducing the deviations between $G-t$ values measured at low vs. high selection function values. The top panel of Fig. 7 demonstrates that most filled circles fall within $1\sigma$ of the global mean $G-t$ value for all cross-matches. This result further justifies our use of the extinction-corrected $G$-band values in subsequent analyses, and it suggests that Galactic dust extinction will not significantly affect our radio-loudness analysis.
3.4 Radio loudness vs. redshift
To perform a further test of systematics, we looked for possible changes in radio loudness as a function of QSO redshift. As shown in Fig. 7 (bottom panel), we observed a moderately significant bias in the average $G-t$ distribution (i.e. in the proxy for radio loudness) at the lowest redshifts ($z<0.5$). We estimated this statistic for the three $G$-bins that we considered above, and we found almost identical trends as for the $G<20.5$ case.
The source of this biased $G-t$ distribution at the low-redshift part of the Quaia QSO catalogue remains unclear. One possibility is that the star–galaxy–QSO separation in the GaiaWISE colour space is less efficient, or the photometric redshifts of the quasars might also be biased at the lowest redshifts where the training set based on Sloan Digital Sky Survey (SDSS) data is less complete. For a detailed explanation, further measurements and statistical probes would be required, which are beyond the scope of this paper.
Figure 9: Radio-loud fraction of Quaia sources colour-coded as a function of the absolute magnitude $G_{\rm abs}$ and the redshift $z$. We observed significant noise due to small-number statistics at the bright-end pixels, leading to strong fluctuations in the radio-loudness fraction.
3.5 Radio-loud fraction across the sky
Looking for possible large-scale spatial patterns on the sky, we created a pixelized map of the radio-loud QSO fraction within each pixel. In Fig. 8, a rather homogeneous distribution is shown across the full observed area, except close to the Galactic plane.
We note that the fluctuations in radio loudness are naturally stronger at low Galactic latitudes, due to sparsity of the QSOs (see the top panel of Fig. 1). For a conservative analysis, future users of our Quaia–VLASS catalogue may apply specific cuts in the Quaia selection function map [^2] Publicly available at https://zenodo.org/records/10403370, offering convenient ways of limiting the analyses to the cleanest and most complete parts of the sky.
3.6 Radio-loud fraction vs. absolute magnitude
Additionally, we tested the possible dependence of the radio-loud fraction on the absolute magnitudes of the quasars. Following Ivezić et al. (2002), we used the absolute magnitude $G_{\rm abs}$, which we calculated as
$G_{\rm abs}=G+5-5\cdot\log\left(d_{\rm lm}(z)\right),$ (5)
where $d_{\rm lm}(z)$ is the luminosity distance [^3] Assuming a Planck $\Lambda$CDM cosmology (Planck Collaboration et al., 2020) of the object, calculated with astropy [^4] https://www.astropy.org/.
We highlight here that we consider this as an exploratory test, because the Quaia photometric redshifts have a non-negligible uncertainty compared to more ideal spectroscopic redshifts of QSOs. These errors also propagate to our estimation of the luminosity distances, and therefore to the $G_{\rm abs}$ values.
Given the above caveats, our main findings regarding the radio-loud fraction patterns in $G_{\rm abs}$ and redshift bins are illustrated in Fig. 9. Each pixel displays the radio-loud fraction of Quaia sources in that bin, using a colour-coding scheme to easily identify any possible trend in the data along either axis.
The figure allows a separate analysis of constant $z$ values in rows against varying $G_{\rm abs}$ values in columns, or vice versa. We found that for a fixed redshift value, radio-loud fraction increases towards brighter $G_{\rm abs}$ bins. Similarly, for a fixed $G_{\rm abs}$, the binned radio-loud fraction values show a decreasing trend with increasing redshift.
We note that, besides the known redshift errors, our $G_{\rm abs}-z$ plot in Fig. 9 displays significant noise due to the sparsity of objects in pixels with the most negative $G_{\rm abs}$ values, leading to large fluctuations in radio-loud fractions. Hence we suggest that QSOs falling in these pixels should be flagged or ignored in any subsequent analyses of redshift- and magnitude-dependence of radio-loud fraction that use our Quaia–VLASS catalogue.
Figure 10: Normalized histograms of the radio-loud and radio-quiet populations as a function of $G_{\rm abs}$. On the left, we show the full sample of cross-matches, while on the right a $G<18.5$ cut is applied.
3.7 Radio-loud vs. radio-quiet source characteristics
In our analysis, we mostly followed Ivezić et al. (2002) and explored if radio-loud and radio-quiet QSOs are statistically consistent with each other, including differently binned subsets. To this end, they compared the normalized histograms of radio-loud and radio-quiet sources as a function of the $i$-band absolute magnitude (similar to our Fig. 10). Above all, they identified an excess of radio-quiet QSOs in the faintest bins, followed by a more abundant radio-loud population, and then again more radio-quiet sources at the peak of the distribution.
Ivezić et al. (2002) attributed these features to noise fluctuations and claimed that differences might only be caused by a few dozen sources. In our analysis, we created similar histograms (see the right panel in Fig. 10) using $G_{\rm abs}$ (limited to objects with $G<18.5$), as opposed to $i_{\rm abs}$ and an $i<18.5$ cut in previous work. Further, we analyzed more sources with 7,431 radio-loud QSOs at $G<18.5$, compared to their 280 radio-loud sources at $i<18.5$. Consequently, our larger data set is expected to help clarify the nature of the excess radio-loud populations in some $G_{\rm abs}$ bins, that were identified as random patterns in previous work (Ivezić et al., 2002).
Considering the brighter $G<18.5$ subset of our Quaia–VLASS quasars, we found that an excess of radio-loud sources does persist at $G_{\rm abs}\approx-24$, and given the statistical errors in our sample, this pattern certainly reaches several standard deviations in three neighbouring bins (see the right panel in Fig. 10).
On the left side of Fig. 10, we present an analogous histogram for the full Quaia–VLASS sample ($G<20.5$) for a more inclusive analysis. We found that the excess of radio-loud sources persists even in the full sample, albeit in a less pronounced manner. In this larger catalogue, we also noted a similar radio-loud excess for the brightest sources at $-30<G_{\rm abs}<-27$. Identifying its origin and exact significance is beyond the scope of this paper, as they might be attributed to yet-unknown systematic effects in the data, or they might be sourced by the Quaia photometric redshift errors.
4 Summary and conclusions
Driven by the availability of modern survey data sets to improve our overall understanding of the astrophysical and cosmological properties of QSOs, we cross-matched the optically selected Quaia (Gaia+WISE) catalogue of quasars with the VLASS radio sources. In our methodology, we closely followed Ivezić et al. (2002) who also analysed the properties of radio-loud QSOs in a similar context, but using significantly smaller data sets. Our main results are the following:
  1. 1.
    We created a Quaia–VLASS catalogue of 43,650 QSOs at declinations $\delta>-40\degr$ with matched radio counterparts, and we make this new catalogue publicly available for further analyses.
  2. 2.
    We measured the radio loudness as a function of the $G$ apparent magnitude, and found consistent results with Ivezić et al. (2002) considering a similar magnitude limit ($G<18.5$).
  3. 3.
    We identified weak trends of radio-loudness changes as a function of the Quaia completeness and as a function of redshift, both pointing to possible systematic effects as sources.
  4. 4.
    We followed up on the findings by Ivezić et al. (2002) in the context of $G_{\rm abs}$ absolute magnitude statistics, and confirmed the previously seen excess of radio-loud quasars at about $G_{\rm abs}\approx-24$, for the brightest optical sources ($G<18.5$).
In summary, here we provided a detailed presentation of the Quaia–VLASS catalogue, and performed various statistical analyses in different subsets of the data to spot possible discrepancies or unexpected trends. While our findings already provided new insights on the radio properties of quasars, the upcoming release of the VLASS Epoch3 catalogue will facilitate an even more advanced analysis. With sources much fainter in the radio, we will further explore the dichotomy between radio-loud and radio-quiet sources, as well as possible relations between the positions of the QSOs in the large-scale cosmic web around them, among other potential applications.
Data availability
The Quaia quasar catalogue and its mask and selection function are publicly available [^5] https://zenodo.org/records/10403370. The VLASS radio source catalogue is also publicly available [^6] https://cirada.ca/vlasscatalogueql0, and we will make our cross-matched Quaia–VLASS data sets, codes and user instructions available publicly [^7] See http://konkoly.hu for information about our data..
Acknowledgments
The Large-Scale Structure (LSS) research group at Konkoly Observatory has been supported by a Lendület excellence grant by the Hungarian Academy of Sciences (MTA). This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement number 101130774. Funding for this project was also available in part through the Hungarian National Research, Development and Innovation Office (NKFIH, grants OTKA NN147550 and K134213). S.F. was supported by the NKFIH excellence grant TKP2021-NKTA-64. L.S.-M was partially supported by the Bulgarian Ministry of Education and Science under Agreement D01-326/04.12.2023. The authors thank Kate Storey-Fisher and the Quaia team for their help with the input quasar catalogue.